Nine cartons of juice cost the same as five fruit cups. Also, one fruit cup costs 50 cents more than one bowl of soup, while one bowl of soup costs 50 cents more than one carton of juice. What would be the cost of each item; a carton of juice, a fruit cup, and a bowl of soup?
Answer by venugopalramana(3286)

Nine cartons of juice cost the same as five fruit cups. Also, one fruit cup costs 50 cents more than one bowl of soup, while one bowl of soup costs 50 cents more than one carton of juice. What would be the cost of each item; a carton of juice, a fruit cup, and a bowl of soup?
let 1 carton of juice cost=x cents
50 cents more =x+50=cost of 1 bowl of soup
50 cents more than above =x+50+50=x+100 = cost of 1 fruit cup
nine cartons of juice costs =9*x cents
five fruit cups cost =5*(x+100)=5x+500 cents
hence 9x=5x+500
9x-5x=500
4x=500
x=500/4=125 cents.
hence 1 carton of juice cost=x =125 cents
cost of 1 bowl of soup =x+50=125+50-175 cents
cost of 1 fruit cup =x+100=125+100=225 cents